OpenStudy (anonymous):

Please help. Write the equation 5x+y=4 in slope-intercept form. A. y=-5x+4 B. y=-5x-4 C. y=5x-4 D. y=5x+4

OpenStudy (anonymous):

subtract 5x from both sides

OpenStudy (anonymous):

so 5x-5x+y=4-5x y=-5x+4 follow? :)